# Fedor V. Fomin

### v2

@@ 1.2 Overview of Theorems 1 and 2
The interplay among the independence number, connectivity, Hamiltonicity, and path covers has long been a popular theme in graph theory, with historical roots going back to the works of Dilworth [Dil50], Gallai and Milgram [GM60], Nash-Williams [NW71], and Chvátal and Erdős [CE72]. Surveys by Jackson and Ordaz [JO90], Bondy [Bon95], and Kawarabayashi [Kaw01] offer excellent overviews of this area. However, most of the research thus far has been centered on establishing combinatorial conditions for Hamiltonicity (or related properties), rather than on algorithmic extensions.
@@
Theorem 1 contributes to a growing line of parameterized algorithms that extend classic theorems from extremal graph theory. Examples include algorithmic adaptations of Poljak–Turzík’s theorem [MPSS14], Dirac’s theorem [FGSS22a], and Hajnal–Szemerédi’s theorem [GHH23]. Such algorithmic extensions can be seen as part of a broader trend in parameterized complexity called above (or below) guarantee parameterization, where one observes that: the natural parameterization of a maximization or minimization problem by the solution size is not satisfactory if there is a lower bound for the solution size that is sufficiently large. The term above guarantee parameterization was introduced by Mahajan and Raman [MRS09], and it has proved fruitful for studying numerous fundamental problems in parameterized complexity and kernelization [AGK+10, BCDF19, CJM+13, FGL+21, FGSS22b, GP16a, GKLM11, GvIMY12, GP16b, GRSY07, LNR+14, GM22]. A feature distinguishing Theorem 1 from the previous work is that it provides “tractability below intractable parameter”.
@@
The exploration of long paths and cycles in graphs represents a well-established and actively pursued research direction within parameterized algorithms. Since the seminal introduction of the color-coding technique by Alon, Yuster, and Zwick [AYZ95], substantial progress in algorithmic methodologies has been achieved. Notably, a range of powerful algorithmic methods has been developed, including those by Bjorklund [Bjö14], Koutis [Kou08], Williams [Wil09], and Zehavi [Zeh16]. This progress has also contributed to our understanding of significantly more general and complex problems, such as topological minor embedding [GKMW11, FLP+20]. A significant portion of parameterized algorithms deals with optimization problems on sparse graphs when specific graph parameters characterize the sparsity and structure of the graph [CNP+11, BCKN13, CKN13, FLPS16, BST23], see also [CFK+15]. Such parameters include treewidth, treedepth, or minimum vertex cover size. On the other hand, for more “general” parameters Hamiltonicity becomes intractable. For example, the problem of deciding graph Hamiltonicity is $\operatorname{{\sf W}}[1]$-hard when parameterized by the clique-width of a graph [FGLS10] and is paraNP parameterized by $\alpha$-treewidth [DMS24].
@@
The existence of a Hamiltonian cycle remains NP-complete on planar graphs [GJT76], circle graphs [Dam89] and for several other generalizations of interval graphs [BB86], for chordal bipartite graphs [Mül96] and for split (and therefore also for chordal) graphs [Gol04], but is solvable in linear time in interval graphs [Kei85] and in convex bipartite graphs [Mül96]. The existence of a Hamiltonian path can be decided in polynomial time for cocomparability graphs [DDKS91] and for circular arc graphs [Dam93].
@@
Many of the above mentioned graph classes are hereditary, i.e., closed in the induced subgraph order, and as such can be described by collections of forbidden induced subgraphs. This has led to carefully examining $H$-free graphs, i.e., graph classes with a single forbidden induced subgraph. For many graphs $H$, the class of $H$-free graphs has nice structural properties (e.g., $P_{4}$-free graphs are exactly the cographs) or their structural properties can be used to design polynomial algorithms for various graph theory problems (as a recent example, cf. colouring of $P_{6}$-free graphs in [CSZ19]). For Hamiltonian-type problems, consider three-vertex graphs $H$. For $H=K_{3}$, both Hamiltonian cycle and Hamiltonian path are NP-complete on triangle-free graphs, since they are NP-complete on bipartite graphs [Mül96]. For $H=3K_{1}$, the edgeless graph on three vertices, both Hamiltonian cycle and Hamiltonian path are polynomial time solvable [DGJ81]. The remaining two graphs, $P_{3}$ and $K_{1}+K_{2}$, are induced subgraphs of $P_{4}$, thus the corresponding class of $H$-free graphs is a subclass of cographs, and as such a subclass of cocomparability graphs, in which the Hamiltonian path and Hamiltonian cycle problems are solvable in polynomial time [DDKS91]. However, a complete characterization of graphs $H$ for which Hamiltonian path or Hamiltonian cycle problems are solvable in polynomial time (and for which they are NP-complete) is not in sight. Until now, the complexity has been open even for the next smallest edgeless graph $4K_{1}$. Theorem 2 answers this complexity question for all edgeless forbidden induced subgraphs. In a companion paper Jedličková and Kratochvíl [JK24] explicitly described situations in which a Hamiltonian path for $\alpha(G)=3,4$ and $5$ in which a Hamiltonian path does exist.
@@
The parameterized complexity of many graph problems remains insufficiently explored for parameters capturing a graph’s density—such as its independence number. It is surprising that, despite the substantial progress in understanding Hamiltonicity, connectivity, and topological minors on graphs of bounded treewidth (which are sparse), relatively little has been achieved for graphs of bounded independence number (which are typically dense). Only a handful of results ($\operatorname{{\sf FPT}}$ or $\operatorname{{\sf X}P}$) address parameterization by a graph’s independence number. These include the $\operatorname{{\sf X}P}$ algorithm of Ovetsky Fradkin and Seymour [FS15] for the edge-disjoint paths problem in directed graphs, the kernelization algorithms of Lochet et al. [LLM+20] for various cycle-hitting problems in directed graphs, the subexponential parameterized algorithms of Misra et al. [MSSZ23] for digraphs with constant independence number, and the recent work of Maalouly, Steiner, and Wulf [MSW23] on $\operatorname{{\sf FPT}}$ algorithms for exact matching.
@@ 2 Preliminaries
We consider simple undirected graphs without loops or multiple edges. The vertex set of a graph $G$ is denoted by $V(G)$, its edge set by $E(G)$. Edges are considered as two-element sets of vertices, thus we write $u\in e$ to express that a vertex $u$ is incident to an edge $e$. For the sake of brevity, we write $uv$ instead of $\{u,v\}$ for the edge containing vertices $u$ and $v$. We say that $u$ is adjacent to $v$ if $uv\in E(G)$. The degree of a vertex is the number of other vertices adjacent to it. The subgraph of $G$ induced by vertices $S\subseteq V(G)$ will be denoted by $G[S]$. We use $G-S$ for $G[V(G)\setminus S]$.
@@
A path in a graph $G$ is a sequence of distinct vertices such that any two consecutive ones are adjacent. The length of a path is the number of its edges. A cycle is formed by a path of length greater than 1 which connects two adjacent vertices. The path (cycle) is Hamiltonian if it contains all vertices of the graph. The path cover number of $G$, denoted by $\mathrm{pc}(G)$, is the smallest number of vertex disjoint paths that cover all vertices of $G$.
@@
A vertex cut in a graph $G$ is a set $A\subset V(G)$ of vertices such that the graph $G-A=G[V(G)\setminus A]$ is disconnected. The vertex connectivity $c_{v}(G)$ of a graph $G$ is the order of a minimum cut in $G$, or $|V(G)|-1$ if $G$ is a complete graph. Since we will not consider edge connectivity, we will often omit the adjective when talking about the connectivity measure, we always have vertex connectivity in mind.
@@
We build upon the following results of Chvátal and Erdős which we will later generalize to Hamiltonian linkages in Theorem 2.
Proposition 1.
[CE72] Let $G$ be an $s$-connected graph.
@@
3 Proof of Theorem 1: Path cover bellow Gallai-Milgram
In this section, we prove Theorem 1. Recall that in the Path Cover problem, we are given a graph $G$ and integer $p\geq 1$, and the task is to decide whether it is possible to cover $G$ with at most $p$ vertex-disjoint paths. Throughout this section, by $\operatorname{pc}(G)$, we denote the minimum number of vertex-disjoint paths required to cover $G$. A well-known theorem by Gallai and Milgram [GM60] guarantees that $\operatorname{pc}(G)\leq\alpha(G)$. When we combine this theorem with Theorem 2, we obtain an algorithm solving Path Cover in time $2^{\alpha(G)^{\mathcal{O}(\alpha(G)^{2})}}\cdot|G|^{\mathcal{O}(1)}$. Our contribution in Theorem 1 makes a step further by establishing the tractability of Path Cover for parameterization $k=\alpha(G)-p$, that is “below” the bound of Gallai-Milgram’s theorem. We restate Theorem 1 here for the reader’s convenience.
@@ 3.1 Small separators: Reducing connected components
The first result is dealing with small vertex separators of $G$. It provides a useful tool that for a small vertex separator $S$ of $G$, reduces the number of connected components in $G-S$ and the value of $\operatorname{pc}(G)$ simultaneously.
@@
Let $G$ be a graph, $S\subseteq V(G)$, and let $C_{1},C_{2},\ldots,C_{t}\subseteq V(G)\setminus S$ be disjoint sets such that for each $i\in[t]$, $C_{i}$ is a clique of $G$ inducing a connected component of $G-S$. Then there is a polynomial time algorithm that, given $G$, $S$, and $\{C_{i}\}_{i\in[t]}$, outputs a set of at most $2|S|^{2}$ indices $X\subseteq[t]$ such that $G$ has a minimum path cover containing a path with vertex set $C_{i}$ for each $i\in[t]\setminus X$.
@@ 3.2 The case of $\operatorname{pc}(G)<k^{C}$
Given $G,k$ and a path cover $\mathcal{P^{\prime}}$ of $G$ of size $p^{\prime}$, the algorithm $\mathcal{A}$ tries to find an optimal path cover of $G$. To achieve that, the algorithm iterates an integer $i$ from $1$ up to $p^{\prime}$ and an integer $j$ from 00 to $i-1$. On each iteration, the algorithm constructs a graph $H_{ij}$ defined as the disjoint union of $j$ copies of $K_{1}$ and $i-j$ copies of $K_{2}$, that is, $G$ contains $j$ isolated vertices and $i-j\geq 1$ edges forming a matching. Notice that $G$ admits a path cover with $i$ paths if and only if $H_{ij}$ has a spanning TM-model in $G$ for some $j\in\{0,\ldots,i-1\}$ because $G$ has at least one edge and, therefore, at least one path in a path cover of minimum size is nontrivial. The algorithm $\mathcal{A}$ thus applies the algorithm of Theorem 2 as a subroutine to $H_{ij},G$ and trivial list assignment $L\equiv V(G)$, with integer parameter set to $p^{\prime}+k=\mathcal{O}(k^{C})$. The subroutine works in
$2^{|H_{ij}|^{\mathcal{O}(p^{\prime}+k)}}\cdot 2^{(p^{\prime}+k)^{\mathcal{O}((p^{\prime}+k)^{2})}}\cdot n^{\mathcal{O}(1)}=2^{k^{\mathcal{O}(k^{2C})}}\cdot n^{\mathcal{O}(1)}$
@@
Note that $\mathcal{A}$ necessary stops and produces a correct solution at some iteration, as it iterates $i$ over $\{1,2,\ldots,p^{\prime}\}$ while $1\leq\operatorname{pc}(G)\leq p^{\prime}$. The proof of the lemma is complete. ∎
@@ 3.3 Proof of Theorem 1
When we reach the situation that the obtained path cover $\mathcal{P}$ and graph $G$ that are irreducible, we deduce that the instance has nice structural properties. In this case, the final solution subroutine calls the algorithms of Lemmas 1 and 2.
The algorithm starts with initializing the path cover $\mathcal{P}$ with $|V(G)|$ paths of length 00. Suppose that at some stage of the algorithm, we have constructed a path cover with $m$ paths for some $m\geq 1$. Let the current set of paths be $\mathcal{P}=\{P_{1},P_{2},\ldots,P_{m}\}$. We denote the endpoints of $P_{i}$ by $s_{i}$ and $t_{i}$.
The first reduction rule is straightforward and comes from the proof of the Gallai-Milgram theorem for undirected graphs.
@@
The correctness of 1 is trivial, and it reduces the number of paths in $\mathcal{P}$ by one. Clearly, the exhaustive application of 1 to $\mathcal{P}$ yields $m\leq\alpha(G)$, as $\{s_{1},s_{2},\ldots,s_{p}\}$ induces an independent set in $G$. When 1 cannot be applied, then the only edges in $G$ allowed between the endpoints of paths in $\mathcal{P}$ are the edges of form $s_{i}t_{i}$.
@@
Proof of Claim 3. Assume that the reduction rules are not applicable, but there are paths $P_{i},P_{j}\in\mathcal{P}$ with $i\neq j$ with an edge between $u_{i}\in V(P_{i})$ and $u_{j}\in V(P_{j})$. If both $P_{i}$ and $P_{j}$ consist of more than two vertices, then 3 is applicable. Thus, without loss of generality, $P_{j}$ consists of one or two vertices, so $V(P_{j})=\{s_{j},t_{j}\}$ and $u_{j}\in\{s_{j},t_{j}\}$. If $P_{i}$ consists of more than two vertices, then 2 is applicable, which is a contradiction. Then, the length of $P_{i}$ is less than three, implying that $u_{i}\in\{s_{i},t_{i}\}$. In this case, 1 is applicable, leading us to the final contradiction. ∎
@@
Proof of Claim 4. We construct an independent set $I$ consisting of more than $m+k$ vertices. Since 1 is not applicable, there are more than $m-k$ special paths. For each special path $P_{i}$ such that $V(P_{i})$ is not a clique, put two independent vertices of $V(P_{i})$ in $I$. For each other special path, put its arbitrary vertex into $I$.
By 3, $I$ is an independent set in $G$. Since there are at least $2k$ non-cliques among vertex sets of special paths, $|I|>(m-k)+2k=m+k$. ∎
@@
The algorithm then exploits the edges between special and usual paths in $\mathcal{P}$. The following reduction rule necessary reduces the number of special paths in $\mathcal{P}$ while not increasing the total number of paths in $\mathcal{P}$.
@@
Figure 2: Illustration of transformations performed in (a) 3 and (b) 4.. Paths before the transformations are horizontal. Paths after the transformations are highlighted in bold and have pairwise-distinct colors. Dashed lines represent absence of an edge. Snake lines represent paths instead of single edges.
@@
If $P_{i}$ and $P_{j}$ are special paths for $i\neq j$, and there is a usual path $P_{\ell}$, and there are two edges $u_{i}x_{i},u_{j}x_{j}$ in $G$ with $u_{i}\in V(P_{i}),u_{j}\in V(P_{j}),x_{i},x_{j}\in V(P_{\ell})$ and $x_{i}$ goes before $x_{j}$ in $P_{\ell}$, then
@@
We first show that $|V(Q_{i})|\geq 3$. Since $\{u_{i},v_{i},x_{i},s_{\ell}\}\subset V(Q_{i})$, it is enough to show that there are at least three distinct vertices among $u_{i},v_{i},x_{i},s_{\ell}$. The only two pairs of vertices that can coincide are $u_{i},v_{i}$ and $x_{i},s_{\ell}$, since they belong to disjoint paths $P_{i}$ and $P_{\ell}$ correspondingly. Hence, if $|V(Q_{i})|<3$, then $u_{i}=v_{i}$ and $x_{i}=s_{\ell}$ both hold. But $u_{i}=v_{i}$ implies that $P_{i}$ has length 00 and $s_{i}=t_{i}=u_{i}$. Then $u_{i}x_{i}=s_{i}s_{\ell}$, so $s_{i}s_{\ell}\in E(G)$, and 1 is applicable for $P_{i}$ and $P_{\ell}$. This contradiction proves $|V(Q_{i})|\geq 3$.
It is left to show that there is no edge between the endpoints of $Q_{i}$, $v_{i}$ and $s_{\ell}$. Indeed, if $v_{i}s_{\ell}\in E(G)$, then we have an edge between a special path $P_{i}$ and an endpoint of $P_{\ell}$, then 2 is applicable. This contradiction proves the claim. ∎
@@
If none of the Reduction Rules 1-4 and Solution Subroutines 1,2 can be applied to $\mathcal{P}$, then there exists $S\subseteq V(G)$ with $|S|<k$ such that $V(P_{i})$ is a connected component in $G-S$ for at least $m-2k$ special paths $P_{i}$. Moreover, such a set $S$ could be constructed in polynomial time.
@@
We show that at least $m-2k$ special paths are isolated in $G-S$. First note that there are at least $m-k$ special paths in $\mathcal{P}$, since the number of usual paths is bounded.
Assume now that $P_{j}$ is a special path that is not isolated in $G-S$. Then there is an edge $uv$ with $u\in V(P_{i})$, $v\in V(P_{j})$, $i\neq j$ and $u\notin S$. By 3, $P_{i}$ can only be usual. Since $u$ is a connector of $P_{i}$ and $u\notin S$, $P_{i}$ has two or more distinct connectors. As 4 is not applicable, all these connectors can only connect $P_{i}$ with $P_{j}$. Hence, each usual path $P_{i}$ gives at most one non-isolated special path in $G-S$. Thus, at most $k$ special paths in $G-S$ are not isolated, implying that at least $(m-k)-k=m-2k$ special paths are isolated in $G-S$. ∎
@@
    Order paths in $\mathcal{P}$ so for each $i\in[h]$, the vertices $V(P_{i})$ induce a complete connected component in $G-S$;

### v3

@@ 1.3 Related Work
The interplay among the independence number, connectivity, Hamiltonicity, and path covers has long been a popular theme in graph theory, with historical roots going back to the works of Dilworth [29], Gallai and Milgram [45], Nash-Williams [82], and Chvátal and Erdős [19]. Surveys by Jackson and Ordaz [59], Bondy [15], and Kawarabayashi [65] offer excellent overviews of this area. However, most of the research thus far has been centered on establishing combinatorial conditions for Hamiltonicity (or related properties), rather than on algorithmic extensions.
@@
Theorem 1 contributes to a growing line of parameterized algorithms that extend classic theorems from extremal graph theory. Examples include algorithmic adaptations of Poljak–Turzík’s theorem [80], Dirac’s theorem [37], and Hajnal–Szemerédi’s theorem [46]. Such algorithmic extensions can be seen as part of a broader trend in parameterized complexity called above (or below) guarantee parameterization, where one observes that: the natural parameterization of a maximization or minimization problem by the solution size is not satisfactory if there is a lower bound for the solution size that is sufficiently large. The term above guarantee parameterization was introduced by Mahajan and Raman [77], and it has proved fruitful for studying numerous fundamental problems in parameterized complexity and kernelization [1, 9, 20, 34, 38, 39, 49, 52, 56, 54, 55, 74, 53]. A feature distinguishing Theorem 1 from the previous work is that it provides “tractability below an intractable parameter”.
@@
The exploration of long paths and cycles in graphs represents a well-established and actively pursued research direction within parameterized algorithms. Since the seminal introduction of the color-coding technique by Alon, Yuster, and Zwick [2], substantial progress in algorithmic methodologies has been achieved. Notably, a range of powerful algorithmic methods has been developed, including those by Björklund [10], Koutis [70], Williams [90], and Zehavi [91]. This progress has also contributed to our understanding of significantly more general and complex problems, such as topological minor embedding [51, 41]. A significant portion of parameterized algorithms deals with optimization problems on sparse graphs when specific graph parameters characterize the sparsity and structure of the graph [23, 13, 22, 40, 4], see also [21]. Such parameters include treewidth, treedepth, or minimum vertex cover size. On the other hand, for more “general” parameters Hamiltonicity becomes intractable. For example, the problem of deciding graph Hamiltonicity is $\operatorname{{\sf W}}[1]$-hard when parameterized by the clique-width of a graph [35] and is $\operatorname{{\sf Para-NP}}$-hard parameterized by $\alpha$-treewidth (see [24] for the definition) as the problem is $\operatorname{{\sf NP}}$-complete on split graphs [50].
@@
The existence of a Hamiltonian cycle remains NP-complete on planar graphs [48], circle graphs [25] and for several other generalizations of interval graphs [7], for chordal bipartite graphs [81] and for split (and therefore also for chordal) graphs [50], but is solvable in linear time in interval graphs [66] and in convex bipartite graphs [81]. The existence of a Hamiltonian path can be decided in polynomial time for cocomparability graphs [27] and for circular arc graphs [26].
@@
Many of the above mentioned graph classes are hereditary, i.e., closed in the induced subgraph order, and as such can be described by collections of forbidden induced subgraphs. This has led to carefully examining $H$-free graphs, i.e., graph classes with a single forbidden induced subgraph. For many graphs $H$, the class of $H$-free graphs has nice structural properties (e.g., $P_{4}$-free graphs are exactly the cographs) or their structural properties can be used to design polynomial algorithms for various graph theory problems (as a recent example, cf. colouring of $P_{6}$-free graphs in [18]). For Hamiltonian-type problems, consider three-vertex graphs $H$. For $H=K_{3}$, both Hamiltonian cycle and Hamiltonian path are NP-complete on triangle-free graphs, since they are NP-complete on bipartite graphs [81]. For $H=3K_{1}$, the edgeless graph on three vertices, both Hamiltonian cycle and Hamiltonian path are polynomial time solvable [31]. The remaining two graphs, $P_{3}$ and $K_{1}+K_{2}$, are induced subgraphs of $P_{4}$; thus, the corresponding class of $H$-free graphs is a subclass of cographs, and as such a subclass of cocomparability graphs, in which the Hamiltonian path and Hamiltonian cycle problems are solvable in polynomial time [27]. However, a complete characterization of graphs $H$ for which Hamiltonian path or Hamiltonian cycle problems are solvable in polynomial time (and for which they are NP-complete) is not in sight. Until now, the complexity has been open even for the next smallest edgeless graph $4K_{1}$. Theorem 2 answers this complexity question for all edgeless forbidden induced subgraphs. In a companion paper Jedličková and Kratochvíl [62] explicitly described situations for $\alpha(G)=3,4$ and $5$ in which a Hamiltonian path does exist.
@@
The parameterized complexity of many graph problems remains insufficiently explored for parameters capturing a graph’s density—such as its independence number. It is surprising that, despite the substantial progress in understanding Hamiltonicity, connectivity, and topological minors on graphs of bounded treewidth (which are sparse), relatively little has been achieved for graphs of bounded independence number (which are typically dense). Only a handful of results ($\operatorname{{\sf FPT}}$ or $\operatorname{{\sf X}P}$) address parameterization by a graph’s independence number. These include the $\operatorname{{\sf X}P}$ algorithm of Ovetsky Fradkin and Seymour [44] for the edge-disjoint paths problem in directed graphs, the kernelization algorithms of Lochet et al. [72] for various cycle-hitting problems in directed graphs, the subexponential parameterized algorithms of Misra et al. [79] for digraphs with constant independence number, and the recent work of Maalouly, Steiner, and Wulf [76] on $\operatorname{{\sf FPT}}$ algorithms for exact matching.
@@ 2 Preliminaries
We consider simple undirected graphs without loops or multiple edges. The vertex set of a graph $G$ is denoted by $V(G)$, its edge set by $E(G)$. Edges are considered as two-element sets of vertices, thus we write $u\in e$ to express that a vertex $u$ is incident to an edge $e$. For the sake of brevity, we write $uv$ instead of $\{u,v\}$ for the edge containing vertices $u$ and $v$. We say that $u$ is adjacent to $v$ if $uv\in E(G)$. The degree of a vertex is the number of other vertices adjacent to it. The subgraph of $G$ induced by vertices $S\subseteq V(G)$ will be denoted by $G[S]$. We use GSG-S for $G[V(G)\setminus S]$.
@@
A path in a graph $G$ is a sequence of distinct vertices such that any two consecutive ones are adjacent. The length of a path is the number of its edges. A cycle is formed by a path of length greater than 1 which connects two adjacent vertices. The path (cycle) is Hamiltonian if it contains all the vertices of the graph. The path cover number of $G$, denoted by $\mathrm{pc}(G)$, is the smallest number of vertex disjoint paths that cover all vertices of $G$.
@@
A vertex cut in a graph $G$ is a set $A\subset V(G)$ of vertices such that the graph $G-A=G[V(G)\setminus A]$ is disconnected. The vertex connectivity $c_{v}(G)$ of a graph $G$ is the order of a minimum cut in $G$, or |V(G)|1|V(G)|-1 if $G$ is a complete graph. Since we will not consider edge connectivity, we will often omit the adjective when talking about the connectivity measure, we always have vertex connectivity in mind.
@@
We build upon the following results of Chvátal and Erdős, which we will later generalize to Hamiltonian linkages in Theorem 2.
Proposition 1 (Chvátal [19]).
Let $G$ be an $s$-connected graph.
@@
3 Proof of Theorem 1: Path cover below Gallai-Milgram
In this section, we prove Theorem 1. Recall that in the Path Cover problem, we are given a graph $G$ and integer $p\geq 1$, and the task is to decide whether it is possible to cover $G$ with at most $p$ vertex-disjoint paths. Throughout this section, by $\operatorname{pc}(G)$, we denote the minimum number of vertex-disjoint paths required to cover $G$. A well-known theorem by Gallai and Milgram [45] guarantees that $\operatorname{pc}(G)\leq\alpha(G)$. When we combine this theorem with Theorem 2, we obtain an algorithm solving Path Cover in time $2^{\alpha(G)^{\mathcal{O}(\alpha(G)^{2})}}\cdot|G|^{\mathcal{O}(1)}$. Our contribution in Theorem 1 makes a step further by establishing the tractability of Path Cover for parameterization $k=\alpha(G)-p$, that is “below” the bound of Gallai-Milgram’s theorem. We restate Theorem 1 here for the reader’s convenience.
@@ 3.1 Small separators: Reducing connected components
The first result is dealing with small vertex separators of $G$. It provides a useful tool that for a small vertex separator $S$ of $G$, reduces the number of connected components in GSG-S and the value of $\operatorname{pc}(G)$ simultaneously.
@@
Let $G$ be a graph, $S\subseteq V(G)$, and let $C_{1},C_{2},\ldots,C_{t}\subseteq V(G)\setminus S$ be disjoint sets such that for each $i\in[t]$, $C_{i}$ is a clique of $G$ inducing a connected component of GSG-S. Then there is a polynomial time algorithm that, given $G$, $S$, and $\{C_{i}\}_{i\in[t]}$, outputs a set of at most $2|S|^{2}$ indices $X\subseteq[t]$ such that $G$ has a minimum path cover containing a path with vertex set $C_{i}$ for each $i\in[t]\setminus X$.
@@ 3.2 The case of $\operatorname{pc}(G)<k^{C}$
Given $G,k$ and a path cover $\mathcal{P^{\prime}}$ of $G$ of size $p^{\prime}$, the algorithm $\mathcal{A}$ tries to find an optimal path cover of $G$. To achieve that, the algorithm iterates an integer $i$ from $1$ up to $p^{\prime}$ and an integer $j$ from 0 to i1i-1. On each iteration, the algorithm constructs a graph $H_{ij}$ defined as the disjoint union of $j$ copies of $K_{1}$ and iji-j copies of $K_{2}$, that is, $H_{ij}$ contains $j$ isolated vertices and $i-j\geq 1$ edges forming a matching. Notice that $G$ admits a path cover with $i$ paths if and only if $H_{ij}$ has a spanning TM-model in $G$ for some $j\in\{0,\ldots,i-1\}$ because $G$ has at least one edge and, therefore, at least one path in a path cover of minimum size is nontrivial. The algorithm $\mathcal{A}$ thus applies the algorithm of Theorem 2 as a subroutine to $H_{ij},G$ and trivial list assignment $L\equiv V(G)$, with integer parameter set to $p^{\prime}+k=\mathcal{O}(k^{C})$. The subroutine works in
$2^{|H_{ij}|^{\mathcal{O}(p^{\prime}+k)}\cdot(p^{\prime}+k)^{\mathcal{O}((p^{\prime}+k)^{2})}}\cdot n^{\mathcal{O}(1)}=2^{k^{\mathcal{O}(k^{2C})}}\cdot n^{\mathcal{O}(1)}$
@@
Note that $\mathcal{A}$ necessarily stops and produces a correct solution at some iteration, as it iterates $i$ over $\{1,2,\ldots,p^{\prime}\}$ while $1\leq\operatorname{pc}(G)\leq p^{\prime}$. The proof of the lemma is complete. ∎
@@ 3.3 Proof of Theorem 1
When we reach the situation where the obtained path cover $\mathcal{P}$ and graph $G$ are irreducible, we deduce that the instance has nice structural properties. In this case, the final solution subroutine calls the algorithms of Lemmas 1 and 2.
The algorithm starts with initializing the path cover $\mathcal{P}$ with $|V(G)|$ paths of length 0. Suppose that at some stage of the algorithm, we have constructed a path cover with $m$ paths for some $m\geq 1$. Let the current set of paths be $\mathcal{P}=\{P_{1},P_{2},\ldots,P_{m}\}$. We denote the endpoints of $P_{i}$ by $s_{i}$ and $t_{i}$.
The first reduction rule is straightforward and comes from the proof of Gallai-Milgram’s theorem for undirected graphs.
@@
The correctness of 1 is trivial, and it reduces the number of paths in $\mathcal{P}$ by one. Clearly, the exhaustive application of 1 to $\mathcal{P}$ yields $m\leq\alpha(G)$, as $\{s_{1},s_{2},\ldots,s_{m}\}$ induces an independent set in $G$. When 1 cannot be applied, then the only edges in $G$ allowed between the endpoints of paths in $\mathcal{P}$ are the edges of form $s_{i}t_{i}$.
@@
Proof of Claim 3. Assume that the reduction rules are not applicable, but there are special paths $P_{i},P_{j}\in\mathcal{P}$ with $i\neq j$ with an edge between $u_{i}\in V(P_{i})$ and $u_{j}\in V(P_{j})$. If both $P_{i}$ and $P_{j}$ consist of more than two vertices, then 3 is applicable. Thus, without loss of generality, $P_{j}$ consists of one or two vertices, so $V(P_{j})=\{s_{j},t_{j}\}$ and $u_{j}\in\{s_{j},t_{j}\}$. If $P_{i}$ consists of more than two vertices, then 2 is applicable, which is a contradiction. Then, the length of $P_{i}$ is less than three, implying that $u_{i}\in\{s_{i},t_{i}\}$. In this case, 1 is applicable, leading us to the final contradiction. ∎
@@
Proof of Claim 4. We construct an independent set $I$ consisting of more than $m+k$ vertices. Since 1 is not applicable, there are more than mkm-k special paths. For each special path $P_{i}$ such that $V(P_{i})$ is not a clique, put two independent vertices of $V(P_{i})$ in $I$. For each other special path, put its arbitrary vertex into $I$.
By 3, $I$ is an independent set in $G$. Since there are at least $2k$ non-cliques among vertex sets of special paths, |I|>(mk)+2k=m+k|I|>(m-k)+2k=m+k. ∎
@@
The algorithm then exploits the edges between special and usual paths in $\mathcal{P}$. The following reduction rule reduces the number of special paths in $\mathcal{P}$ while not increasing the total number of paths in $\mathcal{P}$.
@@
Figure 2: Illustration of transformations performed in (a) 3 and (b) 4. Paths before the transformations are horizontal. Paths after the transformations are highlighted in bold and have pairwise-distinct colors. Dashed lines represent absence of an edge. Snake lines represent paths instead of single edges.
@@
If $P_{i}$ and $P_{j}$ are special paths for $i\neq j$, and there is a usual path $P_{\ell}$, and there are two edges $u_{i}x_{i},u_{j}x_{j}$ in $G$ with $u_{i}\in V(P_{i}),u_{j}\in V(P_{j}),x_{i},x_{j}\in V(P_{\ell})$, $x_{i}\neq x_{j}$, and $x_{i}$ goes before $x_{j}$ in $P_{\ell}$, then
@@
We first show that $|V(Q_{i})|\geq 3$. Since $\{u_{i},v_{i},x_{i},s_{\ell}\}\subseteq V(Q_{i})$, it is enough to show that there are at least three distinct vertices among $u_{i},v_{i},x_{i},s_{\ell}$. The only two pairs of vertices that can coincide are $u_{i},v_{i}$ and $x_{i},s_{\ell}$, since they belong to disjoint paths $P_{i}$ and $P_{\ell}$ correspondingly. Hence, if $|V(Q_{i})|<3$, then $u_{i}=v_{i}$ and $x_{i}=s_{\ell}$ both hold. But $u_{i}=v_{i}$ implies that $P_{i}$ has length 0 and $s_{i}=t_{i}=u_{i}$. Then $u_{i}x_{i}=s_{i}s_{\ell}$, so $s_{i}s_{\ell}\in E(G)$, and 1 is applicable for $P_{i}$ and $P_{\ell}$. This contradiction proves $|V(Q_{i})|\geq 3$.
It is left to show that there is no edge between the endpoints of $Q_{i}$, $v_{i}$ and $s_{\ell}$. Indeed, if $v_{i}s_{\ell}\in E(G)$, then we have an edge between a special path $P_{i}$ and an endpoint of $P_{\ell}$ implying that 2 is applicable. This contradiction proves the claim. ∎
@@
If none of the Reduction Rules 1-4 and Solution Subroutines 1,2 can be applied to $\mathcal{P}$, then there exists $S\subseteq V(G)$ with $|S|<k$ such that $V(P_{i})$ is a connected component in GSG-S for at least m2km-2k special paths $P_{i}$. Moreover, such a set $S$ could be constructed in polynomial time.
@@
We show that at least m2km-2k special paths are isolated in GSG-S. First note that there are at least mkm-k special paths in $\mathcal{P}$, since the number of usual paths is bounded.
Assume now that $P_{j}$ is a special path that is not isolated in GSG-S. Then there is an edge $uv$ with $u\in V(P_{i})$, $v\in V(P_{j})$, $i\neq j$ and $u\notin S$. By 3, $P_{i}$ can only be usual. Since $u$ is a connector of $P_{i}$ and $u\notin S$, $P_{i}$ has two or more distinct connectors. As 4 is not applicable, all these connectors can only connect $P_{i}$ with $P_{j}$. Hence, each usual path $P_{i}$ gives at most one non-isolated special path in GSG-S. Thus, at most $k$ special paths in GSG-S are not isolated, implying that at least (mk)k=m2k(m-k)-k=m-2k special paths are isolated in GSG-S. ∎
@@
    Order paths in $\mathcal{P}$ so for each $i\in[h]$, the vertices $V(P_{i})$ induce a complete connected component in GSG-S;